OpenStudy (anonymous):

helpp with 5t-2-t=14? howw yu got thee answerr

OpenStudy (anonymous):

4t-2=14 (combine like terms) 4t=16 (add 2 to other side) t=4 (divide by 4)

OpenStudy (anonymous):

5t-t = 4t.....14 + 2=16.....16=4t....t=4

OpenStudy (anonymous):

5t-2-t=14 +2 +2 5t - t = 16 4t = 16 /4 /4 t = 4